OpenStudy (hpfan101):

Find the derivative of: y=x^2 e^-3x

OpenStudy (anonymous):

You need to use the product rule here

OpenStudy (anonymous):

\[(f(x)g(x))'=f'(x)g(x)+f(x)g'(x)\]

OpenStudy (anonymous):

just treat f(x)=x^2 and g(x)=e^-3x

OpenStudy (anonymous):

Calculate it and post your answer and I will let you know if you did it correctly

OpenStudy (hpfan101):

Okay so when we apply the power rule would it be like this: (2x)(e^-3x)+(x^2)(e^-3x)

OpenStudy (hpfan101):

Or would the derivative of g(x) be: -3e^x?

OpenStudy (anonymous):

You are almost correct.... except when you take the derivative: \[\frac{d}{dx}e^u=e^u\frac{d}{dx}u=e^uu'\]

OpenStudy (anonymous):

notice how the exponential remains unchanged

OpenStudy (anonymous):

So the first term ((2x)(e^-3x) is correct but the second needs work

OpenStudy (hpfan101):

Oh okay. So for the second term would e^u be x^2 and u would be e^-3x?

OpenStudy (anonymous):

No no e^u was meant to stand in for the exponential factor... the term where you take the derivative of x^2 was the first term (which you did correctly)... I was focusing on how to take the derivative of the exponential factor in the second term properly

OpenStudy (anonymous):

Once you correctly differentiate the exponential factor you multiply it by x^2 (unchanged) and that is the second term in the sum

OpenStudy (hpfan101):

Oh okay. So e^-3x differentiate would be -3e^3x So the second part overall would be x^2(-3e^-3x)

OpenStudy (anonymous):

You missed the minus sign in the first part of that post but your final answer is correct :D now put it all together

OpenStudy (hpfan101):

And then the final answer would be e^-3x(2x-3x^2) since we can factor out e^-3x

OpenStudy (hpfan101):

Ok cool! Thanks for the help! :)

OpenStudy (anonymous):

Correct :D :D you can also factor out an additional x if you want

OpenStudy (anonymous):

Your welcome!